5 ml 1N HCl, 20 ml N/2 H2SO4 and 30 ml N/3 HNO3 are mixed together and the volume is made up to 1L. Normality of the resultant solution is ?

Grade:11

1 Answers

Arun
4 years ago
As we know M(eq) = N x V
where 
M(eq) = equivalent molarity
N = normality of solution
V = volume of solution
Applying this formula for each we get
For HCl----------M(eq) = 1 x 5 = 5M
For H₂SO₄---------M(eq) =  1/2 x 20 = 10M
For HNO₃----------M(eq) = 1/3 x 30 = 10M
For total soluion, M(eq) = M₁ + M₂ + M₃
M(eq) = 5 + 10 + 10 = 25M
and volume = 1000 mL
Applying same formula M(eq) = N x V
25 = N x 1000
N = 25 / 1000
Normality = 0.04 N